DoveDoveDSOve11:

Look at the rectangle and the square: a rectangle pqrs and square lmno are drawn side by side. the length sr of the rectangle is labeled as 12 inches, and the width qr is labeled as 6 inches. the side lm of the square is labeled as 6 inches. sam says that the length of diagonal sq is two times the length of diagonal om. is sam correct? justify your answer and show all your work. your work should state the theorem you used to find the lengths of the diagonals

Hoodmemes:

Can u post the pic of the shapes?

DoveDoveDSOve11:

jhonyy9:

do you can calcule the length of these diagonals ? (rectangle and square ?

jhonyy9:

SQ = ? OM = ?

jhonyy9:

just use Pythagora's theorem

DoveDoveDSOve11:

I did. But the question is asking to show my work and honestly i have no clue how to.

jhonyy9:

how you get the length of SQ ?

jhonyy9:

there is a right triangle for what you need calcule the length of hypotenuse

DoveDoveDSOve11:

I know that i have to half the rectangle, and use the formula for pythagorean theorem, but i dont know how to set it up.

jhonyy9:

ok just make what i ve said please

jhonyy9:

there is the right triangle QRS how you calcule the length of hypotenuse ?

DoveDoveDSOve11:

a^2+ b^2 =c^2......6^2 + 12^2 = c^2.....36 + 144 = c^2.....180 = c^2.....I have to do the square root of 180 and i get 13.41.

jhonyy9:

not is necessary remember the length of QS = sqrt 180 yes ? so and now please calcule the length of OM ?

jhonyy9:

heay ? are you there ?

DoveDoveDSOve11:

Yes Hold on

jhonyy9:

ok please calcule than the length of OM

DoveDoveDSOve11:

Ok, if I half the square i get 

jhonyy9:

sorry what is this ?

jhonyy9:

this is a right triangle with sides length 6 and 6 and you need calcule the length of hypotenuse

jhonyy9:

ok so the length of OM is sqrt 72 yes ?

DoveDoveDSOve11:

me solving it..........................a^2+b^2=c^2...6^2+6^2=c^2.......36+36=c^2.....72=c^2

DoveDoveDSOve11:

Yes, yes it is

jhonyy9:

so you get in the first for QS sqrt180 and for OM sqrt 72 and what you need to prove it now ?

DoveDoveDSOve11:

Im confused, is that the answer.

DoveDoveDSOve11:

The question asks is the kid right?

DoveDoveDSOve11:

or wrong

jhonyy9:

sam says that the length of diagonal sq is two times the length of diagonal om.

jhonyy9:

how you write this ?

DoveDoveDSOve11:

It is

DoveDoveDSOve11:

I have no clue.

jhonyy9:

QS sqrt180 and for OM sqrt 72 sqrt 180 = 2*sqrt72 is this true ? prove it

jhonyy9:

how you prove this ?

DoveDoveDSOve11:

Idk

jhonyy9:

look please what you get than squared both sides ?

DoveDoveDSOve11:

Sorry my bad. Sam is wrong because the square root of 72 is not equal to the square root of 45

jhonyy9:

eliminate the radicals

DoveDoveDSOve11:

How do i do that?

jhonyy9:

sqrt72 ok but how you get this sqrt45 ?

jhonyy9:

sqrt 180 = 2*sqrt72 both sides squared sqrt180 *sqrt180 = (2*sqrt72)*(2*sqrt72)

DoveDoveDSOve11:

Sorry, the double of square root 45 is equal to square root of 180

jhonyy9:

look please what i ve wrote above and make the calcule

jhonyy9:

(sqrt180)^2 = (2sqrt72)^2

jhonyy9:

(sqrt180)^2 = ?

jhonyy9:

you dont know ?

DoveDoveDSOve11:

sorry thats 180

jhonyy9:

yes and (sqrt72)^2 = ?

DoveDoveDSOve11:

72

jhonyy9:

yes and (2sqrt72)^2 = ?

DoveDoveDSOve11:

288

jhonyy9:

yes and now compare please 180 is equal 288 ?

DoveDoveDSOve11:

no

jhonyy9:

so the answer ?

DoveDoveDSOve11:

Sam isn't correct.

jhonyy9:

yes ok now ?

DoveDoveDSOve11:

now what.......

DoveDoveDSOve11:

THERES MORE

jhonyy9:

do you understand it clearly yes ?

jhonyy9:

step by step

jhonyy9:

ok so bye bye

DoveDoveDSOve11:

THANKSSS!!!

jhonyy9:

ok was my pleasure any time

jhonyy9:

yw